OpenStudy (anonymous):

it says: find the distance between the two points. the find the midpoint of the line segment joining the two points (-6,-2) , (-3,5) ?

OpenStudy (anonymous):

slope =rise/run

OpenStudy (anonymous):

distance formula root of (x2-x1)^2+ (yz-y1)^2

OpenStudy (anonymous):

sorry its y2 not yz

OpenStudy (anonymous):

isnt is just adding the xs and ys and dividing by 2 for the midpoint?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

see if you can do it in your head as follows: from -6 to -3 is 3 units. from -2 to 5 is 7 units and pythagoras tell us that \[a^2+b^2=c^2\] the distance is \[\sqrt{3^2+7^2}\]